OpenStudy (anonymous):

44% of US adults 18-34 years think that "Made In America" ads boost sales. Another poll show that 78% of adults in the 18-34 years category use social media online. Suppose that 85% of US adults in the 18-34 years category think that "Made in America" ads boost sales or use social media online. If US adult in the 18-34 years category is randomly selected A: What is the probability that the person thinks that " Made in America" ads boost sales and uses social media online? B: What is the probability that made in America ads boost sales given that the person does not use social media onl

OpenStudy (kropot72):

The US adults can be divided into four categories as follows: A = those that use social media online and do not agree that ads boost sales. B = those that use social media online and agree that ads boost sales. C = those that do not use social media online and do not agree that ads boost sales. D = those that do not use social media online and agree that ads boost sales. A + D = 85% B + D = 44% A + B = 78% By solving the above equations you can find the percentages A, B and D Can you do that?

OpenStudy (anonymous):

Not sure where to start.